# Seat Arrangement Problems under B-utility and W-utility

### v1

@@ 4 Envy-freeness (EFA)
The proof is included in Appendix A. ∎
4.1 1-dimensional preferences
@@
Let the seat graph of a Seat Arrangement instance be a cycle with at least four vertices. Let agents’ preferences be $1$-dimensional, with unique positions, and S-utility. Then for every arrangement $\sigma$ an agent will envy one of its neighbours under $\sigma$.
@@
We will prove the lemma by induction. The base case is an instance of Seat Arrangement $I=(P,F,G,S)$, where $P=\{a,b,c,d\}$, agents’ preferences are derived from a $1$-dimensional space where $a<b<c<d$, and $G$ consist of a cycle with $4$ vertices, labelled $v_{1}$ to $v_{4}$, such that $v_{i}$ is a neighbour of $v_{i+1}$ ($i\mod 4$). For this part of the proof, $(x,y,z,w)$ means that agent $x$ is assigned to $v_{1}$, $y$ to $v_{2}$, $z$ to $v_{3}$, and $w$ to $v_{4}$. Without loss of generality, let us consider every arrangement where $a$ is assigned to $v_{1}$ (so arrangements of the type $(a,x,y,z)$). In arrangement
@@
    $\sigma=(a,b,c,d)$, $a$ envies $b$, because $u_{\sigma^{a\leftrightarrow b}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(b)+f_{a}(d)=u_{\sigma}(a)$
@@
    $\sigma=(a,b,d,c)$, $b$ envies $a$, because $u_{\sigma^{b\leftrightarrow a}}(b)=f_{b}(c)+f_{b}(a)>f_{b}(a)+f_{b}(d)=u_{\sigma}(b)$
@@
    $\sigma=(a,c,b,d)$, $a$ envies $c$, because $u_{\sigma^{a\leftrightarrow c}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(c)+f_{a}(d)=u_{\sigma}(a)$
@@
    $\sigma=(a,c,d,b)$, $b$ envies $a$, because $u_{\sigma^{b\leftrightarrow a}}(b)=f_{b}(c)+f_{b}(a)>f_{b}(a)+f_{b}(d)=u_{\sigma}(b)$
@@
    $\sigma=(a,d,b,c)$, $a$ envies $c$, because $u_{\sigma^{a\leftrightarrow c}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(c)+f_{a}(d)=u_{\sigma}(a)$
@@
    $\sigma=(a,d,c,b)$, $a$ envies $b$, because $u_{\sigma^{a\leftrightarrow b}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(b)+f_{a}(d)=u_{\sigma}(a)$
All other arrangements, where agent $a$ is assigned to $v_{2}$, $v_{3}$, or $v_{4}$, are equivalent to these six.
@@
Take any assignment $\sigma$ on all agents but $z$, on the original cycle with $k$ vertices. By the assumptions, there is at least one agent $b$ who is envious of one of its neighbours $c$. Let $a$ be the other neighbour of $b$, and $d$ the other neighbour of $c$. Let $\sigma^{\prime}$ be an arrangement of the agents onto a cycle with $k+1$ vertices, in which every agent has the same two neighbours as in $\sigma$, except for two neighbouring agents $p$ and $q$, who in $\sigma^{\prime}$ have $z$ between them. That is, we take $\sigma$, and between $p$ and $q$ we introduce a new vertex which is assigned to $z$. If, in $\sigma^{\prime}$, we place $z$ anywhere except between $a$ and $b$, between $b$ and $c$, or between $c$ and $d$, then $b$ will still be envious of $c$, because $u_{\sigma}(b)=u_{\sigma^{\prime}}(b^{\prime})$ and $u_{\sigma^{b\leftrightarrow c}}(b)=u_{\sigma^{\prime b\leftrightarrow c}}(b^{\prime})$. We now look at each of these three cases separately.
@@
Case 1: let $z$ be arranged between $c$ and $d$ in $\sigma^{\prime}$, with no agent envying a neighbour. As $b$ was envious of $c$ and is not anymore, we have that $f_{b}(d)+f_{b}(c)>f_{b}(a)+f_{b}(c)\geq f_{b}(c)+f_{b}(z)$, so $d(b,d)<d(a,b)\leq d(b,z)$. As $c$ is not envious of $b$ in $\sigma^{\prime}$, $f_{c}(a)+f_{c}(b)\leq f_{c}(b)+f_{c}(z)$, from which we get that $d(c,z)\leq d(a,c)$. As $z$ is the rightmost agent, this inequality implies that $c$ is to the right of $a$. Likewise, as $c$ is not envious of $z$, $f_{c}(z)+f_{c}(d)\leq f_{c}(b)+f_{c}(z)$ and thus $d(b,c)\leq d(c,d)$. $z$ is not envious of $c$, so $f_{z}(b)+f_{z}(c)\leq f_{z}(c)+f_{z}(d)$ and thus $d(z,d)\leq d(b,z)$. Hence $d$ is to the right of $b$ in the $1$-dimensional space. Also, because $b$ is closer to $c$ than to $a$, and $c$ is to the right of $a$, $b$ is to the right of $a$. Similarly, as $c$ is to the right of $b$, but $b$ is closer to $c$ than to $d$, $c$ must be between $a$ and $d$.
@@
Therefore, the possible ways in which these agents are ordered in the $1$-dimensional space are $a<c<b<d<z$, and $a<b<c<d<z$.
@@
However, with the first order we get a contradiction, as $d(c,z)\leq d(a,c)$ implies $d(b,z)<d(a,b)$, but we know that $d(a,b)\leq d(b,z)$. Assuming the second order is true, if the cycle has $5$ vertices, $d$ is next to $a$, so $d$ is envious of $z$, as $f_{d}(z)+f_{d}(a)<f_{d}(z)+f_{d}(c)$. Hence, let $l$ be an agent next to $d$ in the seat graph, which must be to the right of $c$ in the $1$-dimensional space, or else $d$ would envy $z$, because if $l$ is to the left of $c$ then $f_{d}(z)+f_{d}(l)<f_{d}(z)+f_{d}(c)$. However, in this case, $z$ is envious of $d$, as $f_{z}(z)+f_{z}(b)<f_{z}(c)+f_{z}(l)$. Thus, in $\sigma^{\prime}$ an agent will always envy a neighbour.

### v2

@@ 4 Envy-freeness (EFA)
We reduce from Partition Into Triangles (PIT). Let $G=(V,E)$ be an instance from PIT with $3n$ vertices ($n\geq 2$), such that every vertex has degree at least $2$ (otherwise no partition exists). We create an instance $I$ of the seat arrangement problem. For every vertex $v_{i}$ we create an agent $p_{i}\in P$. For agents $p_{i},p_{j}\in P$, $f_{p_{i}}(p_{j})=f_{p_{j}}(p_{i})=1$ if $v_{i}$ and $v_{j}$ are neighbours in $G$, otherwise $f_{p_{i}}(p_{j})=f_{p_{j}}(p_{i})=0$. Additionally, for each $p_{i}\in P$, we create two extra agents, $p_{i}^{1}$ and $p_{i}^{2}$. The preferences of these agents are as follows: $f_{p_{i}}(p_{i}^{1})=f_{p_{i}^{1}}(p_{i})=1$, and $f_{p_{i}^{1}}(p_{i}^{2})=f_{p_{i}^{2}}(p_{i}^{1})=1$. Finally, for $1\leq i\leq 3n$, $f_{p_{i}^{2}}(p_{i+1}^{1})=f_{p_{i+1}^{1}}(p_{i}^{2})=1$, where $i+1$ is taken modulo $3n$. For any other pair of agents $a,b\in P$, $f_{a}(b)=f_{b}(a)=0$. That is, every $p_{i}$ gives a symmetrical valuation of $1$ to $p_{j}$ if $v_{i}$ and $v_{j}$ are neighbours, as well as to $p_{i}^{1}$, while $p_{i}^{1}$ gives a symmetrical valuation of $1$ to $p_{i}^{2}$. Finally, every $p_{i}^{2}$ gives a symmetrical valuation of $1$ to the “next” $p_{i+1}^{1}$. All other valuations are 0. The seat graph $H$ contains $n$ 33-cycles and $3n$ 22-cycles, one for each $i$, $1\leq i\leq 3n$. We will use the $p_{i}^{1}$ and $p_{i}^{2}$ agents to force agents of type $p_{i}$ to be arranged into $3$-cycles in any envy-free arrangement.
Assume there is a partition into triangles in $G$. Then we assign to each $3$-cycle in $H$ the three agents from each triangle. For each $1\leq i\leq 3n$ we assign to a $2$-cycle agents $p_{i}^{1}$ and $p_{i}^{2}$. This is an envy-free arrangement, because every agent is getting the maximum utility possible, namely $1$. An example of such an envy-free arrangement is given in Figure 3.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$
$p_{1}$$p_{2}$$p_{3}$$p_{1}^{1}$$p_{1}^{2}$$p_{2}^{1}$$p_{2}^{2}$$p_{3}^{1}$$p_{3}^{2}$
Figure 3: On the left, subgraph of an example PIT instance, with one triangle from the partition highlighted. On the right, part of an envy-free arrangement.
Now, assume there exists an envy-free arrangement $\sigma$ in $I$. Then at least one agent of type $p_{i}^{1}$ is arranged into a $2$-cycle. Assume otherwise. Then all agents of type $p_{i}^{1}$ occupy all vertices in the $3$-cycles, because there are $3n$ agents of type $p_{i}^{1}$ and $n$ 33-cycles. Thus, all agents of type $p_{i}^{2}$ are in $2$-cycles. In this case, however, every agent of type $p_{i}^{1}$ will have utility 0. For each of these $p_{i}^{1}$, her respective $p_{i}^{2}$ will be in a $2$-cycle, so she will feel envy towards the other agent in said $2$-cycle, a contradiction. Thus, at least some $p_{i}^{1}$ is in a $2$-cycle. Fix this $p_{i}^{1}$. As $p_{i}^{2}$ only gives a valuation of $1$ to $p_{i}^{1}$ and $p_{i+1}^{1}$, she must be placed in a $2$-cycle with either of them, because if $p_{i}^{2}$ is placed in a different $2$-cycle or in a $3$-cycle, her utility will be 0 and she will feel envy towards the agent that is placed in the same cycle as $p_{i}^{1}$. Assume first that $p_{i}^{2}$ is in the same cycle as $p_{i}^{1}$ (the other case, where $p_{i}^{2}$ and $p_{i+1}^{1}$ are in the same $2$-cycle, works analogously). Then, as $f_{p_{i-1}^{2}}(p_{i}^{1})=1$, $p_{i-1}^{2}$ must have utility $1$ in $\sigma$, or otherwise she will envy $p_{i}^{1}$. The only other agent that $p_{i-1}^{2}$ gives a valuation of $1$ to is $p_{i-1}^{1}$, so these two agents must be placed in the same $2$-cycle. But that means that $p_{i-2}^{2}$ and $p_{i-2}^{1}$ must be in the same $2$-cycle too. Therefore, by extending this argument, for every pair of agents $p_{j}^{1}$ and $p_{j}^{2}$, they are allocated to the same $2$-cycle under $\sigma$. If, instead, $p_{i}^{2}$ and $p_{i+1}^{1}$ are in the same $2$-cycle, we get that every pair of agents $p_{j}^{2}$ and $p_{j+1}^{1}$ are in the same $2$-cycle. As there are $6n$ agents of type $p_{i}^{1}$ and $p_{i}^{2}$, and $3n$ of these $2$-cycles, it means that no agent of type $p_{i}$ is arranged into a $2$-cycle, and so all are arranged into the $3$-cycles.
Theorem 7.
Envy-Free Arrangement under W-utility is NP-hard even for symmetric and strict preferences and a seat graph where every connected component is a cycle.
Proof.
We reduce from Partition Into Triangles (PIT). Let $G=(V,E)$ be an instance of PIT with $3n$ vertices, such that every vertex has degree at least $2$ (otherwise no partition exists), and maximum degree $4$. We create an instance $I=(P,F,H,W)$ of the seat arrangement problem. For every vertex $v_{i}\in G$, let $V_{i}$ be the set of neighbours of $v_{i}$. Then, for $v_{j},v_{k}\in V_{i}$, we create agent $p_{i,1}^{j,k}\in P$, with $j<k$. This agent represents the possibility of vertex $v_{i}$ being in the same triangle as vertices $v_{j}$ and $v_{k}$ and hence is created only when $v_{i}$, $v_{j}$ and $v_{k}$ are all adjacent to each other. We also create agents $p_{i,2}^{j,k}$ and $p_{i,3}^{j,k}$. Every agent of the form $p_{i,r}^{j,k}$ belongs to subset $P_{i}\subset P$. As $G$ has maximum degree $4$, $v_{i}$ has at most four neighbours in $G$, and as there are six ways of selecting two elements out of a four-element subset, $|P_{i}|\leq 6\cdot 3=18$. If at this point $|P_{i}|<18$, because the degree of $v_{i}$ is less than $4$, we create extra agents until $|P_{i}|=18$. That is, if $|P_{i}|=18-l$ for some $l\geq 1$, we create $l$ additional agents $q_{i}^{1},q_{i}^{2},...,q_{i}^{l}$, which are then added to $P_{i}$. Finally, we create six extra agents, $s_{1}$, $s_{2}$, and $s_{3}$, and $t_{1}$, $t_{2}$, and $t_{3}$, which will be used to “enforce” that an envy-free arrangement corresponds to a partition into triangles. Let $S=\{s_{1},s_{2},s_{3}\}$ and $T=\{t_{1},t_{2},t_{3}\}$.
If vertices $v_{i},v_{j},v_{k}$ are all neighbours of each other, $f_{p_{i,1}^{j,k}}(p_{j,1}^{i,k})=f_{p_{j,1}^{i,k}}(p_{i,1}^{j,k})=3$. Similarly $f_{p_{i,1}^{j,k}}(p_{k,1}^{i,j})=f_{p_{k,1}^{i,j}}(p_{i,1}^{j,k})=3$ and $f_{p_{j,1}^{i,k}}(p_{k,1}^{i,j})=f_{p_{k,1}^{i,j}}(p_{j,1}^{i,k})=3$.
To define valuations for agents within the same $P_{i}$ subset, we need an order relation $\prec_{P_{i}}$. Within each $P_{i}$ subset, $p_{i,r}^{j,k}\prec_{P_{i}}p_{i,c}^{\alpha,\beta}$ if $j<\alpha$, or $j=\alpha$ and $k<\beta$, or $j=\alpha$, $k=\beta$ and $r<c$. Likewise, $p_{i,r}^{j,k}\prec_{P_{i}}q_{i}^{l}$ for any $j,k,r,l$, and $q_{i}^{j}\prec_{P_{i}}q_{i}^{k}$ if $j<k$. Thus, for example, $p_{i,1}^{j,k}\prec_{P_{i}}p_{i,2}^{j,k}\prec_{P_{i}}p_{i,3}^{j,k}\prec_{P_{i}}p_{i,1}^{j,m}\prec_{P_{i}}p_{i,2}^{j,m}\prec_{P_{i}}p_{i,3}^{j,m}\prec_{P_{i}}\dots\prec_{P_{i}}q_{i}^{1}\prec_{P_{i}}\dots\prec_{P_{i}}q_{i}^{l}$. Any subset $S\subseteq P_{i}$ also has the order relation $\prec_{S}$, which is inherited from $P_{i}$.
Let us now define a relation between agents from any subset $S\subseteq P_{i}$. For $a,b\in S$, $a\prec_{S}^{\prime}b$ if $a\prec_{S}b$ and $\nexists c\in S$ such that $a\prec_{S}c\prec_{S}b$. If $a\prec_{S}^{\prime}b$, we say that $a$ is the immediate predecessor of $b$ in $S$, and $b$ is the immediate successor of $a$ in $S$. Finally, to ensure that every agent has an immediate predecessor and an immediate successor, the last element of $S$, according to $\prec_{S}$, is the immediate predecessor in $S$ of the first element according to $\prec_{S}$, and the first element is the immediate successor in $S$ of the last element. This means that we can arrange the agents of $P_{i}$, or any subset $S$, in a circle, such that every agent is adjacent to her immediate predecessor and her immediate successor.
We now define valuations for agents within the same $P_{i}$ subset. If $a\prec_{P_{i}}^{\prime}b$, then $f_{a}(b)=f_{b}(a)=2$. For example, $f_{p_{i,1}^{j,k}}(p_{i,2}^{j,k})=f_{p_{i,2}^{j,k}}(p_{i,1}^{j,k})=2$, $f_{p_{i,2}^{j,k}}(p_{i,3}^{j,k})=f_{p_{i,3}^{j,k}}(p_{i,2}^{j,k})=2$ and $f_{q_{i}^{j}}(q_{i}^{j+1})=f_{q_{i}^{j+1}}(q_{i}^{j})=2$. Finally, because in an envy-free arrangement one agent from $P_{i}$ is assigned to a $3$-cycle, and the remaining agents to the same $17$-cycle (as we are going to see below), we define valuations for some pairs of agents within $P_{i}$. First, let $Q_{i}\subseteq P_{i}$ be a subset of $P_{i}$ that contains every agent from $P_{i}$ except those of the form $p_{i,1}^{j,k}$. We assume that $\prec_{Q_{i}}$ is inherited from $P_{i}$. For every agent of the form $p_{i,2}^{j,k}$, let $a\in Q_{i}$ be her immediate predecessor in $Q_{i}$. This agent $a$ is either of the form $p_{i,3}^{\alpha,\beta}$ or $q_{i}^{l}$, and so is not the immediate predecessor of $p_{i,2}^{j,k}$ in $P_{i}$. Then $f_{p_{i,2}^{j,k}}(a)=f_{a}(p_{i,2}^{j,k})=1.5$. An intuition behind these valuations is that we want agents to give a high valuation towards their immediate predecessor and successor, because those are going to be her neighbours in an envy-free arrangement. Because in an envy-free arrangement one agent of the form $p_{i,1}^{j,k}$ is not assigned to the same cycle as the other agents from $P_{i}$, we need to ensure that some agents within $P_{i}$ give a slightly smaller valuation to agents that are their immediate predecessors or successors after removing said $p_{i,1}^{j,k}$. We refer to Figure 4 for an illustration of this ordering and these valuations.
Let us now define the valuations of agents in $S$ and $T$. For agents $s_{i},s_{j}$, $f_{s_{i}}(s_{j})=f_{s_{j}}(s_{i})=3$, and similarly for agents $t_{i},t_{j}$, $f_{t_{i}}(t_{j})=f_{t_{j}}(t_{i})=3$, where in both cases $1\leq i,j\leq 3$. For agents $s_{i},t_{j}$, $1\leq i,j\leq 3$, $f_{s_{i}}(t_{j})=f_{t_{j}}(s_{i})=2$. Finally, for agents $s_{i},t_{j},a$, $1\leq i,j\leq 3$ and $a\notin S\cup T$, $f_{s_{i}}(a)=f_{a}(s_{i})=1$, and $f_{t_{j}}(a)=f_{a}(t_{j})=-1$. For every other pair of agents $a,b$, $f_{a}(b)=f_{b}(a)=0$. These valuations are not strict, so in order to make them strict we arbitrarily break ties in valuations, in a manner similar to the technique used in the proof of Theorem 5. Thus, to each valuation we add arbitrarily small, positive, and pairwise distinct $\varepsilon_{i}$, where $\varepsilon_{i}$ is defined as in the proof of Theorem 5.
The seat graph $H$ consists of $n+2$ cycles of size $3$, and $3n$ cycles of size $17$, one for each vertex in the PIT instance.
[…]
$v_{1}$$v_{2}$$v_{3}$$v_{4}$
$p_{1,1}^{2,3}$$p_{2,1}^{1,3}$$p_{3,1}^{1,2}$$p_{1,2}^{2,3}$$p_{1,3}^{2,3}$$p_{1,1}^{2,4}$$p_{1,2}^{2,4}$$p_{1,3}^{2,4}$$p_{1,1}^{3,4}$$p_{1,2}^{3,4}$$p_{1,3}^{3,4}$$q_{1}^{1}$$q_{1}^{2}$$q_{1}^{3}$$q_{1}^{4}$$q_{1}^{5}$$q_{1}^{6}$$q_{1}^{7}$$q_{1}^{8}$$q_{1}^{9}$
[…]
An example of part of this arrangement is given in Figure 4. This arrangement, which we call $\sigma$, is valid, because for each $P_{i}$, exactly one agent is assigned to a vertex in a $3$-cycle and the other $17$ agents are arranged into the same $17$-cycle. It is also an envy-free arrangement, as we now show.
Lemma 2.
The arrangement $\sigma$ obtained from the partition into triangles in $G$ is envy-free.
Proof.
Under $\sigma$, agents from $S$ (respectively $T$) have utility at least $3$. If they swap seats with an agent in the same cycle, their utility is unchanged, because they have the same set of neighbours. If they swap seats with an agent from a different graph component, at least one of their neighbours will not be in $S$ (respectively $T$), which means that they will give a valuation strictly smaller than $3$ to that agent, which under W-utility means that their utility is less than $3$.
Similarly, agents belonging to some $P_{i}$ who are arranged into a $3$-cycle have utility at least $3$. If they swap seats with an agent from the same cycle their utility is unchanged, and if they swap seats with any other agent, they will give a valuation smaller than $3$ to at least one of their neighbours, and so their utility would be less than $3$.
Let us now consider agents assigned to a $17$-cycle. For every such agent $a$, either (i) $a$ is arranged next to two agents to which $a$ gives a valuation of at least $2$, or (ii) to an agent to which $a$ gives a valuation of at least $2$ and another agent to which $a$ gives a valuation of at least $1.5$. In case (i), if $a\neq p_{i,1}^{j,k}$ for any $j,k$, then $a$ is adjacent to the two agents that $a$ gives a highest valuation to, which means that if $a$ were to swap seats her utility would decrease. If $a=p_{i,1}^{j,k}$ for some $j,k$, then there are two agents that $a$ gives a valuation of at least $3$ to, which are $p_{j,1}^{i,k}$ and $p_{k,1}^{i,j}$. However, these two agents are in two $17$-cycles such that $a$ gives a valuation of less than $1$ to every other agent in these cycles. Hence, if $a$ were to swap seats her utility would decrease.
Finally, consider case (ii), where $a$ is assigned to a $17$-cycle and gives a valuation of at least $2$ to one neighbour and a valuation between $1.5$ and $2$ to another. In this case $a=q_{i}^{l}$, $a=p_{i,2}^{j,k}$ or $a=p_{i,3}^{j,k}$, for some $l,j,k$, and there exists an agent $p_{i,1}^{\alpha,\beta}$ such that $f_{a}(p_{i,1}^{\alpha,\beta})\geq 2$. However, $p_{i,1}^{\alpha,\beta}$ is assigned to some $3$-cycle, and $a$ gives a valuation of less than $1$ to the other two agents in the $3$-cycle, so her utility would decrease if she swaps seats to be adjacent to $p_{i,1}^{\alpha,\beta}$. Similarly, the utility of $a$ would decrease if she were to swap seats with any other agent, because she would give a valuation of less than $1.5$ to at least one of her neighbours. ∎
Conversely, assume there exists an envy-free arrangement $\sigma$ in $I$. We are going to show that this implies that agents in $S$ belong to one $3$-cycle, and agents in $T$ belong to another one. First, assume that some agent $a\notin S\cup T$ has in her neighbourhood some $t_{i}\in T$. Then $u_{\sigma}(a)<0$. As the seat graph $H$ has more than three connected components, $a$ can switch seats with an agent $b$ allocated to a connected component with no agents from $T$, so $a$ would envy $b$. Hence, no agent from $T$ has in her neighbourhood such an $a$, which means that for any $t_{i}$, her two neighbours are in $S\cup T$. Now, assume that some $t_{i}$ is arranged into a $17$-cycle, and her two neighbours are from $S\cup T$. Then, if either of these neighbours are adjacent to an agent not in $S$ or $T$, said neighbour would envy $t_{i}$. Therefore, the neighbours of $t_{i}$ are adjacent to agents in $S\cup T$ too, which means that there are at least $5$ agents from $S\cup T$ arranged together in the same $17$-cycle. As $|S\cup T|=6$, there are either $5$ or all $6$ agents from $S\cup T$ arranged together in the same $17$-cycle. However, this implies that two agents from $S\cup T$ have a neighbour not in $S\cup T$, and thus they envy $t_{i}$, a contradiction to the envy-freeness of $\sigma$.
Hence, all agents in $T$ are arranged into $3$-cycles that only include agents in $S\cup T$. Because $|S\cup T|=6$, either all agents in $T$ belong to the same cycle, or some $t_{i}$ is in a cycle with two agents from $S$, and the other two agents from $T$ are in a cycle with some $s_{j}$. However, in this case $t_{i}$ would envy $s_{j}$. Thus, all agents from $T$ belong to the same $3$-cycle. Now, if some $s_{i}\in S$ has in her neighbourhood an agent $a\notin S\cup T$, then $u_{\sigma}(s_{i})<2$, and thus she will feel envy towards any agent $t_{j}\in T$, because $u_{\sigma^{s_{i}\leftrightarrow t_{j}}}(s_{i})\geq 2$. Hence, all agents in $S$ are allocated to the same $3$-cycle.
We now show that in $\sigma$, the agents assigned to each $17$-cycle belong to the same subset $P_{i}\subset P$. For a contradiction, assume the opposite. Then there are two agents, $a\in P_{i}$ and $b\in P_{j}$, that belong to different subsets and are adjacent to each other. However, because $f_{a}(b)=f_{b}(a)<1$, $u_{\sigma}(a)<1$. As $u_{\sigma^{a\leftrightarrow s_{1}}}(a)\geq 1$, $a$ envies $s_{1}$, a contradiction to the envy-freeness of $\sigma$. Therefore, all agents arranged into a $17$-cycle belong to the same $P_{i}$. As there are $3n$ 1717-cycles and $3n$ $P_{i}$’s, for every $P_{i}$, seventeen agents belonging to it are assigned to some $17$-cycle.
For every $P_{i}$, $|P_{i}|=18$, so exactly one agent from every $P_{i}$ is not assigned to a vertex from a $17$-cycle, and thus is assigned to a $3$-cycle. If this agent is not of type $p_{i,1}^{j,k}$, then she gives a valuation of less than $1$ to every agent outside of $S\cup P_{i}$, and so her utility in $\sigma$ is 0, which means that she envies $s_{1}$, a contradiction. Hence, the only agent from $P_{i}$ not assigned to a $17$-cycle is of type $p_{i,1}^{j,k}$. In order for her utility to be at least $1$ and thus not be envious of $s_{1}$, the two other vertices in the cycle she has been assigned to must be occupied by $p_{j,1}^{i,k}$ and $p_{k,1}^{i,j}$, all of which have utility at least $3$ under $\sigma$. That is, we have a triangle in the original instance with vertices $v_{i},v_{j},v_{k}$. As only one agent from $P_{i}$ is in some cycle component, we form only one triangle with vertex $v_{i}$, meaning that we obtain a partition into triangles in $G$. ∎
@@
Let the seat graph of a seat arrangement instance with S-utility be a cycle with at least four vertices. Let agents’ preferences be $1$-dimensional, with unique positions. Then for every arrangement $\sigma$ an agent will envy one of its neighbours under $\sigma$.
@@
We will prove the lemma by induction. The base case is an instance of seat arrangement $I=(P,F,G,S)$, where $P=\{p,q,r,s\}$, agents’ preferences are derived from a $1$-dimensional space where $l_{p}<l_{q}<l_{r}<l_{s}$, and $G$ consists of a cycle with $4$ vertices, labelled $v_{1}$ to $v_{4}$, such that $v_{i}$ is a neighbour of $v_{i+1}$ ($i\mod 4$). For this part of the proof, $(x,y,z,w)$ means that agent $x$ is assigned to $v_{1}$, $y$ to $v_{2}$, $z$ to $v_{3}$, and $w$ to $v_{4}$. Without loss of generality, let us consider every arrangement where $p$ is assigned to $v_{1}$ (so arrangements of the type $(a,x,y,z)$). In arrangement
@@
    $\sigma=(p,q,r,s)$, $p$ envies $q$, because $u_{\sigma^{p\leftrightarrow q}}(p)=f_{p}(q)+f_{p}(r)>f_{p}(q)+f_{p}(s)=u_{\sigma}(p)$
@@
    $\sigma=(p,q,s,r)$, $q$ envies $p$, because $u_{\sigma^{q\leftrightarrow p}}(q)=f_{q}(r)+f_{q}(p)>f_{q}(p)+f_{q}(s)=u_{\sigma}(q)$
@@
    $\sigma=(p,r,q,s)$, $p$ envies $r$, because $u_{\sigma^{p\leftrightarrow r}}(p)=f_{p}(q)+f_{p}(r)>f_{p}(r)+f_{p}(s)=u_{\sigma}(p)$
@@
    $\sigma=(p,r,s,q)$, $q$ envies $p$, because $u_{\sigma^{q\leftrightarrow p}}(q)=f_{q}(r)+f_{q}(p)>f_{q}(p)+f_{q}(s)=u_{\sigma}(q)$
@@
    $\sigma=(p,s,q,r)$, $p$ envies $r$, because $u_{\sigma^{p\leftrightarrow r}}(p)=f_{p}(q)+f_{p}(r)>f_{p}(r)+f_{p}(s)=u_{\sigma}(p)$
@@
    $\sigma=(p,s,r,q)$, $p$ envies $q$, because $u_{\sigma^{p\leftrightarrow q}}(p)=f_{p}(q)+f_{p}(r)>f_{p}(q)+f_{p}(s)=u_{\sigma}(p)$
All other arrangements, where agent $p$ is assigned to $v_{2}$, $v_{3}$, or $v_{4}$, are equivalent to these six.
@@
Take any assignment $\sigma$ on all agents but $z$, on the original cycle with $k$ vertices. By the assumptions, there is at least one agent $q$ who is envious of one of her neighbours $r$. Let $p$ be the other neighbour of $q$, and $s$ the other neighbour of $r$. Let $\sigma^{\prime}$ be an arrangement of the agents onto a cycle with $k+1$ vertices, in which every agent has the same two neighbours as in $\sigma$, except for two neighbouring agents $a$ and $b$, who in $\sigma^{\prime}$ have $z$ between them. That is, we take $\sigma$, and between $a$ and $b$ we introduce a new vertex which is assigned to $z$. If, in $\sigma^{\prime}$, we place $z$ anywhere except between $p$ and $q$, between $q$ and $r$, or between $r$ and $s$, then $q$ will still be envious of $r$, because $u_{\sigma}(q)=u_{\sigma^{\prime}}(q)$, which is smaller than $u_{\sigma^{q\leftrightarrow r}}(q)=u_{\sigma^{\prime q\leftrightarrow r}}(q)$. We now look at each of these three cases separately.
@@
Case 1: let $z$ be arranged between $r$ and $s$ in $\sigma^{\prime}$, with no agent envying a neighbour. The positions of the agents in $G$ are shown is Figure 5. As $q$ was envious of $r$ and is not anymore, we have that $f_{q}(s)+f_{q}(r)>f_{q}(p)+f_{q}(r)\geq f_{q}(r)+f_{q}(z)$, so $d(q,s)<d(p,q)\leq d(q,z)$. As $r$ is not envious of $q$ in $\sigma^{\prime}$, $f_{r}(p)+f_{r}(q)\leq f_{r}(q)+f_{r}(z)$, from which we get that $d(r,z)\leq d(p,r)$. As $z$ is the rightmost agent, this inequality implies that $r$ is to the right of $p$. Likewise, as $r$ is not envious of $z$, $f_{r}(z)+f_{r}(s)\leq f_{r}(q)+f_{r}(z)$ and thus $d(q,r)\leq d(r,s)$, so $r$ is closer to $q$ than to $s$. $z$ is not envious of $r$, so $f_{z}(q)+f_{z}(r)\leq f_{z}(r)+f_{z}(s)$ and thus $d(z,s)\leq d(q,z)$. Hence $s$ is to the right of $q$ in the $1$-dimensional space. Also, because $q$ is closer to $s$ than to $p$, and $s$ is to the right of $q$, $q$ is to the right of $p$. Finally, $r$ is between $p$ and $s$, because $r$ is to the right of $p$, $s$ is to the right of $q$, and $r$ is closer to $q$ than to $s$, implying than $r$ is to the left of $s$.
@@
Therefore, the possible ways in which these agents are ordered in the $1$-dimensional space are $l_{p}<l_{r}<l_{q}<l_{s}<l_{z}$, and $l_{p}<l_{q}<l_{r}<l_{s}<l_{z}$.
@@
However, with the first order we get a contradiction, as $d(r,z)\leq d(p,r)$ implies $d(q,z)<d(p,q)$, but we know that $d(p,q)\leq d(q,z)$. Assuming the second order is true, if the cycle has $5$ vertices, $s$ is next to $p$, so $s$ is envious of $z$, as $f_{s}(z)+f_{s}(p)<f_{s}(z)+f_{s}(r)$. Hence, let $l$ be an agent next to $s$ in the seat graph. If $l$ is to the left of $r$ in the $1$-dimensional space then $f_{d}(z)+f_{d}(l)<f_{d}(z)+f_{d}(r)$, so $s$ would envy $z$. Thus, $l$ is to the right of $r$. However, in this case, $z$ envies $s$, as $f_{z}(r)+f_{z}(s)<f_{z}(s)+f_{z}(l)$. Thus, in $\sigma^{\prime}$ an agent will always envy a neighbour.
ppqqrrzzss
Figure 5: Positions of the agents in part of the seat graph $G$. Case 11.